OpenStudy (anonymous):

What is 3.2k-4.3=12.6k+14.5

OpenStudy (anonymous):

get K by itself. 3.2k-4.3=12.6k+14.5 3.2k-4.3+4.3=12.6k+14.5+4.3 3.2k=12.6k+18.8 3.2k-12.6k=12.6k-12.6k+18.8 -9.4k=18.8 -9.4k/-9.4=18.8/-9.4 k= -2 Ok??